Solve dy/dx + yx = y2e^(x^2/2)sinx


1
Expert's answer
2021-01-19T18:23:00-0500

"\frac{dy}{dx}+yx=y^2e^{\left(\frac{x^2}{2}\right)}\sin \left(x\right)"


First order Bernoulli Ordinary Differential Equation

A first order Bernoulli ODE has the form of "y'+p\left(x\right)y=q\left(x\right)y^n"

Substitute "\frac{dy}{dx}\mathrm{\:with\:}y'\:"

"y'\:+yx=y^2e^{\left(\frac{x^2}{2}\right)}\sin \left(x\right)"

Rewrite in the form of a first order Bernoulli ODE:  "y'+p\left(x\right)y=q\left(x\right)y^n"

"p\left(x\right)=x,\:\quad q\left(x\right)=e^{\frac{x^2}{2}}\sin \left(x\right),\:\quad n=2"

The general solution is obtained by substituting "v=y^{1-n}"

and solving "\frac{1}{1-n}v'+p\left(x\right)v=q\left(x\right)"

Transform to  "\frac{1}{1-n}v'+p\left(x\right)v=q\left(x\right):"         "-v'+xv = e^{\left(\frac{x^2}{2}\right)}\sin \left(x\right)"

Solve "-v'+xv = e^{\left(\frac{x^2}{2}\right)}\sin \left(x\right)" :                  "v = e^{\large\frac{x^2}{2}}cos(x) + c_1 e^{\large\frac{x^2}{2}}"

Substitute back "v = y {-1}:"       "y^{-1} = e^{\large\frac{x^2}{2}}cos(x) + c_1 e^{\large\frac{x^2}{2}}"

Isolate y :          "y = \large\frac{1}{e^{\frac{x^2}{2}}\left(\cos \left(x\right)+c_1\right)}"


The answer:     "y = \large\frac{1}{e^{\frac{x^2}{2}}\left(\cos \left(x\right)+c_1\right)}"